The longtable package^{∗}

David Carlisle^{†}

1998/05/13

Abstract

This package defines the longtable environment, a multi-page version of tabular.

List of Tables

1 An optional table caption (used in the list of tables) . . . . . . . . 2 2 A floating table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3 A difficult \multicolumn combination: pass 1 . . . . . . . . . . . . 6 4 A difficult \multicolumn combination: pass 2 . . . . . . . . . . . . 6 5 A difficult \multicolumn combination: pass 3 . . . . . . . . . . . . 6 6 A difficult \multicolumn combination: pass 4 . . . . . . . . . . . . 6 7 A summary of longtable commands . . . . . . . . . . . . . . . . . . 9

1 Introduction

longtable The longtable package defines a new environment, longtable, which has most of the features of the tabular environment, but produces tables which may be broken by TEX’s standard page-breaking algorithm. It also shares some features with the table environment. In particular it uses the same counter, table, and has a similar \caption command. Also, the standard \listoftables command lists tables produced by either the table or longtable environments.

The following example uses most of the features of the longtable environment. An edited listing of the input for this example appears in Section 8.

Note: Various parts of the following table will not line up correctly until this document has been run through L^{A}TEX several times. This is a characteristic feature of this package, as described below.

∗This file has version number v4.09, last revised 1998/05/13.

†The new algorithm for aligning ‘chunks’ of a table used in version 4 of this package was devised coded and documented by David Kastrup, dak@neuroinformatik.ruhr-uni-bochum.de.

. . . . . . . . . . . . . . . . . . . . . . . longtable.sty . . . . . . . . . . . . . . . . . . . . . . .

Table 1: A long table

*  This part appears at the top of the table  *
*  First Second  *
*  longtable columns are specified in the  *
*  same way as in the tabular environment. *

*  @{*}r||p{1in}@{*} in this case.  *

*  Each row ends with a \\ command. *
*  The \\ command has an optional  *
*  argument, just as in the  *
*  tabular environment. *
*  See the effect of \\[10pt] ?  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Also \hline may be used, as in tabular. *
*  That was a \hline .  *

*  That was \hline\hline .  *

              This is a \multicolumn{2}{||c||}

*  If a page break occurs at a \hline then a line is drawn *
*  at the bottom of one page and at the top of the next. *
*  The [t] [b] [c] argument of tabular can not be used.*
*  The optional argument may be one of [l] [r] [c] *
*  to specify whether the table should be adjusted  *
*  to the left, right or centrally.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  This goes at the bottom.  *

. . . . . . . . . . . . . . . . . . . . . . . . . . . Page 2 . . . . . . . . . . . . . . . . . . . . . . . . . .

. . . . . . . . . . . . . . . . . . . . . . . longtable.sty . . . . . . . . . . . . . . . . . . . . . . .

Table 1: (continued)

*  This part appears at the top of every other page  *
*  First Second  *
*Some lines may take up a lot of space, like this: This last column is a “p” column so this “row” of the table can take up several lines. Note however that TEX will never break a page within such a row. Page breaks only occur between rows of the table or at \hline commands.*

*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  Lots^{1} of lines like this.  *
*  Lots of lines like this^{2}  *
*  Lots of lines like this.  *
*  Lots of lines like this.  *
*  These lines will appear  *
*  in place of the usual foot  *
*  at the end of the table  *

2 Chunk Size

LTchunksize In order to TEX multi-page tables, it is necessary to break up the table into smaller chunks, so that TEX does not have to keep everything in memory at one time. By default longtable uses 20 rows per chunk, but this can be set by the user, with e.g., \setcounter{LTchunksize}{10}.^{3} These chunks do not affect page breaking, thus if you are using a TEX with a lot of memory, you can set LTchunksize to be several pages of the table. TEX will run faster with a large LTchunksize.

1This is a footnote.

2longtable takes special precautions, so that footnotes may also be used in ‘p’ columns.

3You can also use the plain TEX syntax \LTchunksize=10.

A  tabular  environment
within a floating table

Table 2: A floating table

However, if necessary, longtable can work with LTchunksize set to 1, in which case the memory taken up is negligible. Note that if you use the commands for setting the table head or foot (see below), the LTchunksize must be at least as large as the number of rows in each of the head or foot sections.

This document specifies \setcounter{LTchunksize}{10}. If you look at the previous table, after the first run of L^{A}TEX you will see that various parts of the table do not line up. L^{A}TEX will also have printed a warning that the column widths had changed. longtable writes information onto the .aux file, so that it can line up the different chunks. Prior to version 4 of this package, this information was not used unless a \setlongtables command was issued, however, now the information is always used, using a new algorithm^{4} and so \setlongtables is no longer needed. It is defined (but does nothing) for the benefit of old documents that use it.

3 Captions and Headings

At the start of the table one may specify lines which are to appear at the top

\endhead of every page (under the headline, but before the other lines of the table). The

lines are entered as normal, but the last \\ command is replaced by a \endhead

\endfirsthead command. If the first page should have a different heading, then this should be entered in the same way, and terminated with the \endfirsthead command. The LTchunksize should be at least as large as the number of rows in the heading.

\endfoot There are also \endfoot and \endlastfoot commands which are used in the same \endlastfoot way (at the start of the table) to specify rows (or an \hline) to appear at the

bottom of each page. In certain situations, you may want to place lines which logically belong in the table body at the end of the firsthead, or the beginning of the lastfoot. This helps to control which lines appear on the first and last page of the table.

\caption The \caption{...} command is essentially equivalent to

\multicolumn{n}{c}{\parbox{\LTcapwidth}{...}}

where n is the number of columns of the table. You may set the width of the caption with a command such as \setlength{\LTcapwidth}{2in} in the preamble of your document. The default is 4in. \caption also writes the information to produce an entry in the list of tables. As with the \caption command in the figure and table environments, an optional argument specifies the text to appear in the list of tables if this is different from the text to appear in the caption. Thus the caption for table 1 was specified as \caption[An optional table caption

(used in the list of tables)]{A long table\label{long}}.

You may wish the caption on later pages to be different to that on the first page. In this case put the \caption command in the first heading, and put a subsidiary caption in a \caption[] command in the main heading. If the optional argument to \caption is empty, no entry is made in the list of tables. Alternatively, if you do not want the table number to be printed each time, use the \caption* command.

4Due to David Kastrup.

The captions are set based on the code for the article class. If you have redefined the standard \@makecaption command to produce a different format for the captions, you may need to make similar changes to the longtable version, \LT@makecaption. See the code section for more details.

A more convenient method of customising captions is given by the caption(2) package, which provides commands for customising captions, and arranges that the captions in standard environments, and many environments provided by packages (including longtable) are modified in a compatible manner.

You may use the \label command so that you can cross reference longtables with \ref. Note however, that the \label command should not be used in a heading that may appear more than once. Place it either in the firsthead, or in the body of the table. It should not be the first command in any entry.

4 Multicolumn entries

The \multicolumn command may be used in longtable in exactly the same way as for tabular. So you may want to skip this section, which is rather technical, however coping with \multicolumn is one of the main problems for an environment such as longtable. The main effect that a user will see is that certain combinations of \multicolumn entries will result in a document needing more runs of L^{A}TEX before the various ‘chunks’ of a table align.

The examples in this section are set with LTchunksize set to the minimum value of one, to demonstrate the effects when \multicolumn entries occur in different chunks.

Consider Table 3. In the second chunk, longtable sees the wide multicolumn entry. At this point it thinks that the first two columns are very narrow. All the width of the multicolumn entry is assumed to be in the third column. (This is a ‘feature’ of TEX’s primitive \halign command.) longtable then passes the information that there is a wide third column to the later chunks, with the result that the first pass over the table is too wide.

If the ‘saved row’ from this first pass was re-inserted into the table on the next pass, the table would line up in two passes, but would be much two wide.

kill The solution to this problem used in Versions 1 and 2, was to use a \kill line. If a line is \killed, by using \kill rather than \\ at the end of the line, it is used in calculating column widths, but removed from the final table. Thus entering \killed copies of the last two rows before the wide multicolumn entry would mean that \halign ‘saw’ the wide entries in the first two columns, and so would not widen the third column by so much to make room for the multicolumn entry.

In Version 3, a new solution was introduced. If the saved row in the .aux file was not being used, longtable used a special ‘draft’ form of \multicolumn, this modified the definition, so the spanning entry was never considered to be wider than the columns it spanned. So after the first pass, the .aux file stored the widest normal entry for each column, no column was widened due to \spanned columns. By default longtable ignored the .aux file, and so each run of L^{A}TEX was considered a first pass. Once the \setlongtables declaration was given, the saved row in the .aux file, and the proper definition of \multicolumn were

. . . . . . . . . . . . . . . . . . . . . . . longtable.sty . . . . . . . . . . . . . . . . . . . . . . .

Table 3: A difficult \multicolumn combination: pass 1

1 2 3

wide multicolumn spanning 1–3

multicolumn 1–2  3
wide 1 2 3

Table 4: A difficult \multicolumn combination: pass 2

1 2 3

wide multicolumn spanning 1–3

multicolumn 1–2  3
wide 1 2 3

Table 5: A difficult \multicolumn combination: pass 3

1 2 3

wide multicolumn spanning 1–3

multicolumn 1–2  3
wide 1 2 3

Table 6: A difficult \multicolumn combination: pass 4

1 2 3

wide multicolumn spanning 1–3

multicolumn 1–2  3
wide 1 2 3

. . . . . . . . . . . . . . . . . . . . . . . . . . . Page 6 . . . . . . . . . . . . . . . . . . . . . . . . . . used. If any \multicolumn entry caused one of the columns to be widened, this information could not be passed back to earlier chunks, and so the table would not correctly line up until the third pass. This algorithm always converged in three passes as described above, but in examples such as the ones in Tables 3–6, the final widths were not optimal as the width of column 2, which is determined by a \multicolumn entry was not known when the final width for column 3 was fixed, due to the fact that both \multicolumn commands were switched from ‘draft’ mode to ‘normal’ mode at the same time.

Version 4 alleviates the problem considerably. The first pass of the table will indeed have the third column much too wide. However, on the next pass longtable will notice the error and reduce the column width accordingly. If this has to propagate to chunks before the \multicolumn one, an additional pass will, of course, be needed. It is possible to construct tables where this rippling up of the correct widths taks several passes to ‘converge’ and produce a table with all chunks aligned. However in order to need many passes one needs to construct a table with many overlapping \multicolumn entries, all being wider than the natural widths of the columns they span, and all occuring in different chunks. In the typical case the algorithm will converge after three or four passes, and, the benefits of not needing to edit the document before the final run to add \setlongtables, and the better choice of final column widths in the case of multiple \multicolumn entries will hopefully more than pay for the extra passes that may possibly be needed.

So Table 3 converges after 4 passes, as seen in Table 6.

You can still speed the convergence by introducing judicious \kill lines, if you happen to have constellations like the above.

If you object even to L^{A}TEX-ing a file twice, you should make the first line of every longtable a \kill line that contains the widest entry to be used in each column. All chunks will then line up on the first pass.

5 Adjustment

The optional argument of longtable controls the horizontal alignment of the table. The possible options are [c], [r] and [l], for centring, right and left adjustment,

\LTleft respectively. Normally centring is the default, but this document specifies

\LTright
 \setlength\LTleft\parindent
 \setlength\LTright\fill

in the preamble, which means that the tables are set flush left, but indented by the usual paragraph indentation. Any lengths can be specified for these two parameters, but at least one of them should be a rubber length so that it fills up the width of the page, unless rubber lengths are added between the columns using the \extracolsep command. For instance

  \begin{tabular*}{\textwidth}{@{\extracolsep{...}}...}

produces a full width table, to get a similar effect with longtable specify

 \setlength\LTleft{0pt}
 \setlength\LTright{0pt}
 \begin{longtable}{@{\extracolsep{...}}...}

6 Changes

This section highlights the major changes since version 2. A more detailed change log may be produced at the end of the code listing if the ltxdoc.cfg file specifies

\AtBeginDocument{\RecordChanges}
\AtEndDocument{\PrintChanges}

Changes made between versions 2 and 3.

• The mechanism for adding the head and foot of the table has been completely rewritten. With this new mechanism, longtable does not need to issue a \clearpage at the start of the table, and so the table may start half way down a page. Also the \endlastfoot command which could not safely be implemented under the old scheme, has been added.

• longtable now issues an error if started in the scope of \twocolumn, or the multicols environment.

• The separate documentation file longtable.tex has been merged with the package file, longtable.dtx using Mittelbach’s doc package.

• Support for footnotes has been added. Note however that \footnote will not work in the ‘head’ or ‘foot’ sections of the table. In order to put a footnote in those sections (e.g., inside a caption), use \footnotemark at that point, and \footnotetext anywhere in the table body that will fall on the same page.

• The treatment of \multicolumn has changed, making \kill lines unnecessary, at the price of sometimes requiring a third pass through L^{A}TEX.

• The \newpage command now works inside a longtable.

Changes made between versions 3 and 4.

• A new algorithm is used for aligning chunks. As well as the widest width in each column, longtable remembers which chunk produced this maximum. This allows it to check that the maximum is still achieved in later runs. As longtable can now deal with columns shrinking as the file is edited, the \setlongtables system is no longer needed and is disabled.

• An extra benefit of the new algorithm’s ability to deal with ‘shrinking’ columns is that it can give better (narrower) column widths in the case of overlapping \multicolumn entries in different chunks than the previous algorithm produced.

• The ‘draft’ multicolumn system has been removed, along with related com-

mands such as \LTmulticolumn.

• The disadvantage of the new algorithm is that it can take more passes. The theoretical maximum is approximately twice the length of a ‘chain’ of columns with overlapping \multicolumn entries, although in practice it usually converges as fast as the old version. (Which always converged in three passes once \setlongtables was activated.)

• \\* and \nopagebreak commands may be used to control page breaking.

7 Summary

Table 7: A summary of longtable commands

Parameters

\LTleft  Glue to the left of the table.  (\fill)
\LTright  Glue to the right of the table.  (\fill)
\LTpre  Glue before the the table.  (\bigskipamount)
\LTpost  Glue after the the table.  (\bigskipamount)
\LTcapwidth  The width of a parbox containing the caption.  (4in)
LTchunksize  The number of rows per chunk.  (20)

Optional arguments to \begin{longtable}

none  Position as specified by \LTleft and \LTright.
[c]  Centre the table.
[l]  Place the table flush left.
[r]  Place the table flush right.

Commands to end table rows

\\  Specifies the end of a row
\\[⟨dim⟩]  Ends row, then adds vertical space (as in the tabular environment).
\\*  The same as \\ but disallows a page break after the row.

\tabularnewline Alternative to \\ for use in the scope of \raggedright and similar

commands that redefine \\.

\kill  Row is ‘killed’, but is used in calculating widths.
\endhead  Specifies rows to appear at the top of every page.
\endfirsthead  Specifies rows to appear at the top the first page.
\endfoot  Specifies rows to appear at the bottom of every page.
\endlastfoot  Specifies rows to appear at the bottom of the last page.

longtable caption commands

\caption{⟨caption⟩} Caption ‘Table ?: ⟨caption⟩’, and a ‘⟨caption⟩’ entry in the list of

tables.

\caption[⟨lot⟩]{⟨caption⟩} Caption ‘Table ?: ⟨caption⟩’, and a ‘⟨lot⟩’ entry in the list of

tables.

\caption[]{⟨caption⟩}  Caption ‘Table ?: ⟨caption⟩’, but no entry in the list of tables.
\caption*{⟨caption⟩}  Caption ‘⟨caption⟩’, but no entry in the list of tables.

Commands available at the start of a row

\pagebreak  Force a page break.
\pagebreak[⟨val⟩]  A ‘hint’ between 0 and 4 of the desirability of a break.
\nopagebreak  Prohibit a page break.
\nopagebreak[⟨val⟩]  A ‘hint’ between 0 and 4 of the undesirability of a break.
\newpage  Force a page break.

Footnote commands available inside longtable

\footnote  Footnotes, but may not be used in the table head & foot.
\footnotemark  Footnotemark, may be used in the table head & foot.
\footnotetext  Footnote text, use in the table body.

Setlongtables

\setlongtables  Obsolete command. Does nothing now.

8 Verbatim highlights from Table 1

\begin{longtable}{@{*}r||p{1in}@{*}}
KILLED & LINE!!!! \kill
\caption[An optional table caption ...]{A long table\label{long}}\\
\hline\hline
\multicolumn{2}{@{*}c@{*}}%
     {This part appears at the top of the table}\\
\textsc{First}&\textsc{Second}\\
\hline\hline
\endfirsthead
\caption[]{(continued)}\\
\hline\hline
\multicolumn{2}{@{*}c@{*}}%
      {This part appears at the top of every other page}\\
\textbf{First}&\textbf{Second}\\
\hline\hline
\endhead
\hline
This goes at the&bottom.\\
\hline
\endfoot
\hline
These lines will&appear\\
in place of the & usual foot\\
at the end& of the table\\
\hline
\endlastfoot
\env{longtable}  columns  are specified& in the \\
same way as  in the \env{tabular}& environment.\\
...
\multicolumn{2}{||c||}{This is a ...}\\
...
Some lines may take...&
    \raggedleft This last column is a ``p'' column...
    \tabularnewline
...
Lots of lines& like this.\\
...
\hline
Lots\footnote{...} of lines& like this.\\
Lots   of   lines& like this\footnote{...}\\
\hline
Lots of lines& like this.\\
...
\end{longtable}

9 The Macros

1 ⟨∗package⟩

9.1 Initial code

Before declaring the package options, we must define some defaults here.

\LT@err The error generating command

 2 \def\LT@err{\PackageError{longtable}}

\LT@warn The warning generating command

 3 \def\LT@warn{\PackageWarning{longtable}}

\LT@final@warn If any longtables have not aligned, generate a warning at the end of the run at

\AtEndDocument.
 4 \def\LT@final@warn{%
 5 \AtEndDocument{%
 6  \LT@warn{Table \@width s have changed. Rerun LaTeX.\@gobbletwo}}%
 7 \global\let\LT@final@warn\relax}

9.2 Options

The first two options deal with error handling. They are compatible with the options used by the tracefnt package.

errorshow Only show errors on the terminal. ‘warnings’ are just sent to the log file.

 8 \DeclareOption{errorshow}{%
 9 \def\LT@warn{\PackageInfo{longtable}}}

pausing Make every warning message into an error so TEX stops. May be useful for debugging.

10 \DeclareOption{pausing}{%
11 \def\LT@warn#1{%
12  \LT@err{#1}{This is not really an error}}}

set final The next options are just alternative syntax for the \setlongtables declaration.

13 \DeclareOption{set}{}
14 \DeclareOption{final}{}
15 \ProcessOptions

9.3 User Setable Parameters

\LTleft \LTright Glue to the left and right of the table, default \fill (ie centred).

16 \newskip\LTleft       \LTleft=\fill
17 \newskip\LTright      \LTright=\fill

\LTpre \LTpost Glue before and after the longtable. \bigskip by default.

18 \newskip\LTpre        \LTpre=\bigskipamount
19 \newskip\LTpost       \LTpost=\bigskipamount

\LTchunksize Chunk size (The number of rows taken per \halign). Default 20.

20 \newcount\LTchunksize \LTchunksize=20
\c@LTchunksize Added in V3.07 to allow the L^{A}TEX syntax \setcounter{LTchunksize}{10}.
21 \let\c@LTchunksize\LTchunksize

\LTcapwidth Width of the \parbox containing the caption. Default 4in.

22 \newdimen\LTcapwidth  \LTcapwidth=4in

9.4 Internal Parameters

\LT@head \LT@firsthead \LT@foot \LT@lastfoot Boxes for the table head and foot.
23 \newbox\LT@head
24 \newbox\LT@firsthead
25 \newbox\LT@foot
26 \newbox\LT@lastfoot

\LT@cols Counter for number of columns.

27 \newcount\LT@cols

\LT@rows Counter for rows up to chunksize.

28 \newcount\LT@rows

\c@LT@tables Counter for the tables, added in V3.02. Previous versions just used the L^{A}TEX counter table, but this fails if table is reset during a document, eg report class resets it every chapter.

This was changed from \newcount\LT@tables in V3.04. L^{A}TEX counters are preserved correctly when \includeonly is used. In the rest of the file \LT@tables has been replaced by \c@LT@tables without further comment.

29 \newcounter{LT@tables}

\c@LT@chunks We need to count through the chunks of our tables from Version 4 on.

30 \newcounter{LT@chunks}[LT@tables]

\c@table \fnum@table \tablename If the table counter is not defined (eg in letter style), define it. (Added in V3.06.)

31 \ifx\c@table\undefined
32 \newcounter{table}
33 \def\fnum@table{\tablename~\thetable}
34 \fi
35 \ifx\tablename\undefined
36 \def\tablename{Table}
37 \fi

\LT@out In a normal style, longtable uses the .aux file to record the column widths. With

letter.sty, use a separate .lta file. (Added in V3.06.)

Not needed for new letter class.

\ifx\startlabels\undefined
  \let\@auxout\@auxout
\else
  {\@input{\jobname.lta}}%

  \newwrite\@auxout
  \immediate\openout\@auxout=\jobname.lta
\fi

\LT@p@ftn Temporary storage for footnote text in a ‘p’ column.

38 \newtoks\LT@p@ftn

\LT@end@pen Special penalty for the end of the table. Done this way to save using up a count register.

39 \mathchardef\LT@end@pen=30000

9.5 The longtable environment

\longtable Called by \begin{longtable}. This implementation does not work in multiple column formats. \par added at V3.04.

40 \def\longtable{%
41 \par
42 \ifx\multicols\@undefined
43 \else
44  \ifnum\col@number>\@ne
45  \@twocolumntrue
46  \fi
47 \fi
48 \if@twocolumn
49  \LT@err{longtable not in 1-column mode}\@ehc
50 \fi
51 \begingroup

Check for an optional argument.

52 \@ifnextchar[\LT@array{\LT@array[x]}}

\LT@array Start setting the alignment. Based on \@array from the L^{A}TEX kernel and the

array package.

Since Version 3.02, longtable has used the internal counter \c@LT@tables. The L^{A}TEX counter table is still incremented so that \caption works correctly.

53 \def\LT@array[#1]#2{%
54 \refstepcounter{table}\stepcounter{LT@tables}%

Set up the glue around the table if an optional argument given.

55 \if l#1%
56  \LTleft\z@ \LTright\fill
57 \else\if r#1%
58  \LTleft\fill \LTright\z@
59 \else\if c#1%
60  \LTleft\fill \LTright\fill
61 \fi\fi\fi

Set up these internal commands for longtable.

  \global\let\LT@mcw@rn\relax
62 \let\LT@mcol\multicolumn

Now redefine \@tabarray to restore \hline and \multicolumn so that arrays and tabulars nested in longtable (or in page headings on longtable pages) work out OK. Saving the original definitions done here so that you can load the array package before or after longtable.

63 \let\LT@@tabarray\@tabarray
64 \let\LT@@hl\hline
65 \def\@tabarray{%
66  \let\hline\LT@@hl
    \let\multicolumn\LT@mcol
67  \LT@@tabarray}%
68 \let\\\LT@tabularcr\let\tabularnewline\\%
69 \def\newpage{\noalign{\break}}%

More or less standard definitions, but first start a \noalign.

70 \def\pagebreak{\noalign{\ifnum`}=0\fi\@testopt{\LT@no@pgbk-}4}%
71 \def\nopagebreak{\noalign{\ifnum`}=0\fi\@testopt\LT@no@pgbk4}%
72 \let\hline\LT@hline \let\kill\LT@kill\let\caption\LT@caption
73 \@tempdima\ht\strutbox
74 \let\@endpbox\LT@endpbox

Set up internal commands according to Lamport or Mittelbach.

75 \ifx\extrarowheight\@undefined

Initialise these commands as in tabular from the L^{A}TEX kernel.

76  \let\@acol\@tabacol
77  \let\@classz\@tabclassz \let\@classiv\@tabclassiv
78  \def\@startpbox{\vtop\LT@startpbox}%
79  \let\@@startpbox\@startpbox
80  \let\@@endpbox\@endpbox
81  \let\LT@LL@FM@cr\@tabularcr
82 \else

Initialise these commands as in array. \d@llar replaced by \d@llarbegin \d@llarend in V3.03 to match array V2.0h. We do not need to set \d@llarbegin and \d@llarend as the array package gives them the correct values at the top level.

83  \advance\@tempdima\extrarowheight
84  \col@sep\tabcolsep
85  \let\@startpbox\LT@startpbox\let\LT@LL@FM@cr\@arraycr
86 \fi

The rest of this macro is mainly based on array package, but should work for the standard tabular too.

87 \setbox\@arstrutbox\hbox{\vrule
88  \@height \arraystretch \@tempdima
89  \@depth \arraystretch \dp \strutbox
90  \@width \z@}%
91 \let\@sharp##\let\protect\relax

Interpret the preamble argument.

92 \begingroup
93  \@mkpream{#2}%

We need to rename \@preamble here as F.M.’s scheme uses \global, and we may need to nest \@mkpream, eg for \multicolumn or an array. We do not need to worry about nested longtables though!

94  \xdef\LT@bchunk{%
95  \global\advance\c@LT@chunks\@ne
96  \global\LT@rows\z@\setbox\z@\vbox\bgroup

The following line was added in v4.05. In order to get the \penalties to work at chunk boundaries Need to take more care about where and when \lineskip glue is added. The following does nothing at top of table, and in header chunks, but in normal body chunks it sets \prevdepth (to 0pt, but any value would do) so that \lineskip glue will be added. the important thing to note is that the glue will be added after any vertical material coming from \noalign.

97  \LT@setprevdepth
98  \tabskip\LTleft\halign to\hsize\bgroup
99  \tabskip\z@ \@arstrut \@preamble \tabskip\LTright \cr}%
100 \endgroup

Find out how many columns we have (store in \LT@cols).

101 \expandafter\LT@nofcols\LT@bchunk&\LT@nofcols

Get the saved row from \LT@i… \LT@ix (from the .aux file), or make a new blank row.

102 \LT@make@row

A few more internal commands for longtable.

103 \m@th\let\par\@empty
104 \everycr{}\lineskip\z@\baselineskip\z@

Start the first chunk.

105 \LT@bchunk}

\LT@no@pgbk Can simplify the standard \@no@pgbk as this is vmode only but then need to close

the \noalign.
106 \def\LT@no@pgbk#1[#2]{\penalty #1\@getpen{#2}\ifnum`{=0\fi}}

\LT@start This macro starts the process of putting the table on the current page. It is not called until either a \\ or \endlongtable command ends a chunk, as we do not know until that point which of the four possible head or foot sections have been specified.

It begins by redefining itself, so that the table is only started once! Until V3.04, was redefined to \relax, now use \endgraf to force the page-breaker to wake up.

107 \def\LT@start{%
108 \let\LT@start\endgraf
109 \endgraf\penalty\z@\vskip\LTpre

Start a new page if there is not enough room for the table head, foot, and one extra line.

110 \dimen@\pagetotal
111 \advance\dimen@ \ht\ifvoid\LT@firsthead\LT@head\else\LT@firsthead\fi
112 \advance\dimen@ \dp\ifvoid\LT@firsthead\LT@head\else\LT@firsthead\fi
113 \advance\dimen@ \ht\LT@foot

At this point I used to add \ht\@arstrutbox and \dp\@arstrutbox as a measure of a row size. However this can fail spectacularly for p columns which might be much larger. Previous versions could end up with the table starting with a foot, then a page break then a head then a ‘first head’! So now measure the first line of the table accurately by \vsplitting it out of the first chunk.

114 \dimen@ii\vfuzz
115 \vfuzz\maxdimen
116  \setbox\tw@\copy\z@
117  \setbox\tw@\vsplit\tw@ to \ht\@arstrutbox
118  \setbox\tw@\vbox{\unvbox\tw@}%
119 \vfuzz\dimen@ii
120 \advance\dimen@ \ht
121  \ifdim\ht\@arstrutbox>\ht\tw@\@arstrutbox\else\tw@\fi
122 \advance\dimen@\dp
123  \ifdim\dp\@arstrutbox>\dp\tw@\@arstrutbox\else\tw@\fi
124 \advance\dimen@ -\pagegoal
125 \ifdim \dimen@>\z@\vfil\break\fi

Store height of page minus table foot in \@colroom.

126  \global\@colroom\@colht

If the foot is non empty, reduce the \vsize and \@colroom accordingly.

127 \ifvoid\LT@foot\else
128  \advance\vsize-\ht\LT@foot
129  \global\advance\@colroom-\ht\LT@foot
130  \dimen@\pagegoal\advance\dimen@-\ht\LT@foot\pagegoal\dimen@
131  \maxdepth\z@
132 \fi

Put the table head on the page, and then switch to the new output routine.

133 \ifvoid\LT@firsthead\copy\LT@head\else\box\LT@firsthead\fi
134 \output{\LT@output}}
\endlongtable Called by \end{longtable}.
135 \def\endlongtable{%

Essentially add a final \\. But as we now know the number of actual chunks, we first strip away all entries referring to a maximum entry beyond the table (this can only happen if a table has been shortened, or the table numbering has gone awry). In that case we at least start collecting valid new information with the last chunk of this table, by removing the width constraint.

136 \crcr
137 \noalign{%
138  \let\LT@entry\LT@entry@chop
139  \xdef\LT@save@row{\LT@save@row}}%
140 \LT@echunk
141 \LT@start
142 \unvbox\z@
143 \LT@get@widths

Write the dummy row to the .aux file. Since V3.06, use .lta for letter.sty.

144 \if@filesw
145  {\let\LT@entry\LT@entry@write\immediate\write\@auxout{%

Since Version 3.02, longtable has used the internal counter \c@LT@tables rather than the L^{A}TEX counter table. This information looks entirely different from version 3 information. Still, we don’t need to rename the macro name because later code will consider the information to have no columns, and thus will throw the old data away.

146  \gdef\expandafter\noexpand
147  \csname LT@\romannumeral\c@LT@tables\endcsname
148  {\LT@save@row}}}%
149 \fi

At this point used to issue a warning if a \multicolumn has been set in draft mode.

  \LT@mcw@rn

If the last chunk has different widths than the first, warn the user. Also trigger a warning to rerun L^{A}TEX at the end of the document.

150 \ifx\LT@save@row\LT@@save@row
151 \else
152  \LT@warn{Column \@width s have changed\MessageBreak
153  in table \thetable}%
154  \LT@final@warn
155 \fi

Force one more go with the longtable output routine.

156 \endgraf\penalty -\LT@end@pen

Now close the group to return to the standard routine.

157 \endgroup

Reset \@mparbottom to allow marginpars close to the end of the table.^{5}

158 \global\@mparbottom\z@
159 \pagegoal\vsize
160 \endgraf\penalty\z@\addvspace\LTpost

Footnotes. As done in the multicol package.

161 \ifvoid\footins\else\insert\footins{}\fi}

9.6 Counting Columns

Columns are counted by examining \@preamble, rather than simply getting \@mkpream to increment the counter as it builds the preamble so that this package works with many of the packages which add extra column specifiers to L^{A}TEX’s standard ones.

Version 1 counted \@sharp’s to calculate the number of columns, this was changed for Version 2 as it does not work with the NFSS. Now count &’s. (lfonts.new (and now the Standard L^{A}TEX definition) defines \@tabclassz so that \@sharp is inside a group.)

\LT@nofcols Find the next &, then look ahead to see what is next.

162 \def\LT@nofcols#1&{%
163 \futurelet\@let@token\LT@n@fcols}

5This can not be the correct. However if it is omitted, there is a problem with marginpars, for example on page 3 of this document. Any Output Routine Gurus out there?

\LT@n@fcols Add one, then stop at an \LT@nofcols or look for the next &. The \expandafter trick was added in Version 3, also the name changed from \@LT@nofcols to preserve the \LT@ naming convention.

164 \def\LT@n@fcols{%
165 \advance\LT@cols\@ne
166 \ifx\@let@token\LT@nofcols
167  \expandafter\@gobble
168 \else
169  \expandafter\LT@nofcols
170 \fi}

9.7 The \\ and \kill Commands

\LT@tabularcr The internal definition of \\. In the * form, insert a \nobreak after the next \cr

(or \crcr).

This star form processing was finally added in v4.05. For the previous six or seven years the comment at this point said

This definition also accepts \\*, which acts in the same way as \\. tabular does this, but longtable probably ought to make \\* prevent page breaking.

{\ifnum0=`}\fi added in version 3.01, required if the first entry is empty. The above in fact is not good enough, as with array package it can introduce a {} group in math mode, which changes the spacing. So use the following variant. Added in v3.14.

171 \def\LT@tabularcr{%
172 \relax\iffalse{\fi\ifnum0=`}\fi
173 \@ifstar
174  {\def\crcr{\LT@crcr\noalign{\nobreak}}\let\cr\crcr
175  \LT@t@bularcr}%
176  {\LT@t@bularcr}}
\LT@crcr
177 \let\LT@crcr\crcr

\LT@setprevdepth This will be redefined to set the \prevdepth at the start of a chunk.

178 \let\LT@setprevdepth\relax
\LT@t@bularcr
179 \def\LT@t@bularcr{%

Increment the counter, and do tabular’s \\ or finish the chunk. The \expandafter trick was added in Version 3. Set the \prevdepth at the start of a new chunk. (Done here so not set in header chunks).

180 \global\advance\LT@rows\@ne
181 \ifnum\LT@rows=\LTchunksize
182  \gdef\LT@setprevdepth{%
183  \prevdepth\z@\global
184  \global\let\LT@setprevdepth\relax}%
185  \expandafter\LT@xtabularcr
186 \else
187  \ifnum0=`{}\fi
188  \expandafter\LT@LL@FM@cr
189 \fi}

\LT@xtabularcr This justs looks for an optional argument.

190 \def\LT@xtabularcr{%
191 \@ifnextchar[\LT@argtabularcr\LT@ntabularcr}

\LT@ntabularcr The version with no optional argument. \ifnum0=`{\fi} added in version 3.01. Changed in 3.14.

192 \def\LT@ntabularcr{%
193 \ifnum0=`{}\fi
194 \LT@echunk
195 \LT@start
196 \unvbox\z@
197 \LT@get@widths
198 \LT@bchunk}

\LT@argtabularcr The version with an optional argument. \ifnum0=`{\fi} added in version 3.01. Changed in 3.14.

199 \def\LT@argtabularcr[#1]{%
200 \ifnum0=`{}\fi
201 \ifdim #1>\z@
202  \unskip\@xargarraycr{#1}%
203 \else
204  \@yargarraycr{#1}%
205 \fi

Add the dummy row, and finish the \halign.

206 \LT@echunk
207 \LT@start
208 \unvbox\z@
209 \LT@get@widths
210 \LT@bchunk}

\LT@echunk This ends the current chunk, and removes the dummy row.

211 \def\LT@echunk{%
212 \crcr\LT@save@row\cr\egroup
213 \global\setbox\@ne\lastbox

The following line was added in v4.05. longtable relies on \lineskip glue (which is 0pt) to provide break points between each row so the table may be split into pages.

Previous releases left the \lineskip glue at the end of each chunk that had been added when the dummy row was added. There was no glue at the start of the next chunk as TEX normally dooes not put \lineskip glue at the top of a box. This meant that normally the chunks fitted together perfectly, however \noalign material at a chunk boundary came before the first row of the next chunk but after the lineskip glue at the end of this chunk. This is the wrong place, e.g., it means even a \penalty10000 does not stop a break as the \lineskip glue in the previous item on the list provides a legal breakpoint. So now remove the \lineskip glue that was before the dummy row and introduce \LT@setprevdepth to set the \prevdepth at the start of the next chunk, to make sure \lineskip glue is added later.

214  \unskip
215 \egroup}

\LT@entry We here give the ‘basic’ definition of \LT@entry, namely that used in alignment templates. It has a \kern only if the maximum is imposed from a different chunk. The \ifhmode test reveals the first entry, when we don’t want to add an &.

216 \def\LT@entry#1#2{%
217 \ifhmode\@firstofone{&}\fi\omit
218 \ifnum#1=\c@LT@chunks
219 \else
220  \kern#2\relax
221 \fi}

\LT@entry@chop This definition for the argument of \LT@save@row is used to scrap all those maxima which could not be verified because they occur after the end of the table. This can happen only if a table has been shortened (or the sequencing got mixed up) since the previous run. Note that this is premature: the last chunk still is going to be set, and with the chopped limits.

222 \def\LT@entry@chop#1#2{%
223 \noexpand\LT@entry
224  {\ifnum#1>\c@LT@chunks
225  1}{0pt%
226  \else
227  #1}{#2%
228  \fi}}

\LT@entry@write To write an entry for the aux file, we use a slightly surprising definition which has the sole purpose of avoiding overfull lines (which might break TEX’s limits when reading the aux file, probably you’d need to have a few hundred columns before this happened but…).

229 \def\LT@entry@write{%
230 \noexpand\LT@entry^^J%
231 \@spaces}

\LT@kill This ends the current chunk as above, but strips off two rows, the ‘dummy row’ and the ‘killed row’ before starting the next chunk. Since V3.04, the old chunk is reboxed at the start of the box containing the next chunk. This allows \kill to be used in headers, which must be processed in a single box.

232 \def\LT@kill{%
233 \LT@echunk
234 \LT@get@widths
235 \expandafter\LT@rebox\LT@bchunk}

\LT@rebox Drop the old chunk (box0) back at the top of the new chunk, removing the killed row. This macro added at V3.04.

236 \def\LT@rebox#1\bgroup{%
237 #1\bgroup
238 \unvbox\z@
239 \unskip
240 \setbox\z@\lastbox}

9.8 The Dummy Row

The dummy row is kept inside of the macro \LT@save@row.

\LT@blank@row \LT@build@blank Create a blank row if we are not using the info in the .aux file.

241 \def\LT@blank@row{%
242 \xdef\LT@save@row{\expandafter\LT@build@blank
243  \romannumeral\number\LT@cols 001 }}

Whoops! What’s that supposed to be? A drop-in replacement for the first task of Appendix D in the TEXbook. The \romannumeral produces \LT@cols instances of m followed by i. The below macro then replaces the ms by appropriate entries.

244 \def\LT@build@blank#1{%
245 \if#1m%
246  \noexpand\LT@entry{1}{0pt}%
247  \expandafter\LT@build@blank
248 \fi}

\LT@make@row Prior to version 4, by default did not use information in the .aux file but now we can define \LT@make@row to use the .aux file, even on the ‘draft’ passes.

249 \def\LT@make@row{%
250 \global\expandafter\let\expandafter\LT@save@row
251  \csname LT@\romannumeral\c@LT@tables\endcsname
252 \ifx\LT@save@row\relax
253  \LT@blank@row

Now a slightly difficult part comes. Before we decide making the template from the .aux file info we check that the number of fields has remained the same. If it hasn’t, either the table format has changed, or we have the wrong table altogether. In both cases, we decide to better drop all gathered information and start over.

The expansion between !… ! below will be empty if the number of \LT@entry macros including arguments in \LT@save@row is equal to \LT@cols. If it is not empty, we throw the row away and start from scratch.

254 \else
255  {\let\LT@entry\or
256  \if!%
257  \ifcase\expandafter\expandafter\expandafter\LT@cols
258  \expandafter\@gobble\LT@save@row
259  \or
260  \else
261  \relax
262  \fi

263 !%

264  \else
265  \aftergroup\LT@blank@row
266  \fi}%
267 \fi}

\setlongtables Redefine \LT@make@row to use information in the .aux file, if there is a saved row

for this table with the right number of columns.

Since Version 3.02, longtable has used the internal counter \c@LT@tables rather than the L^{A}TEX counter table. The warning message was added at V3.04, as was the \global, to stop save-stack overflow.

Since Version 4.01 \setlongtables does nothing as it is not needed, but is defined as \relax for the benefit of old documents.

268 \let\setlongtables\relax

\LT@get@widths This is the heart of longtable. If it were not for the table head and foot, this macro together with the modified \\ command would form the basis of quite a simple little package file for long tables. It is closely modelled on the \endvrulealign macro of appendix D of the TEXbook.

269 \def\LT@get@widths{%

\global added at V3.04, to stop save-stack overflow.

Loop through the last row, discarding glue, and saving box widths. At V3.04 changed the scratch box to 2, as the new \kill requires that \box0 be preserved.

270 \setbox\tw@\hbox{%
271  \unhbox\@ne
272  \let\LT@old@row\LT@save@row
273  \global\let\LT@save@row\@empty
274  \count@\LT@cols
275  \loop
276  \unskip
277  \setbox\tw@\lastbox
278  \ifhbox\tw@
279  \LT@def@row
280  \advance\count@\m@ne
281  \repeat}%

Remember the widths if we are in the first chunk.

282 \ifx\LT@@save@row\@undefined
283  \let\LT@@save@row\LT@save@row
284 \fi}

\LT@def@row Add a column to the dummy row. Name changed from \defLT@save@row in Version 3, to preserve the \LT@ naming convention.

285 \def\LT@def@row{%

We start by picking the respective entry from our old row. These redefinitions of \LT@entry are local to the group started in \LT@get@widths.

286 \let\LT@entry\or
287 \edef\@tempa{%
288  \ifcase\expandafter\count@\LT@old@row
289  \else
290  {1}{0pt}%
291  \fi}%

Now we tack the right combination in front of \LT@save@row:

292 \let\LT@entry\relax
293 \xdef\LT@save@row{%
294  \LT@entry
295  \expandafter\LT@max@sel\@tempa
296  \LT@save@row}}

\LT@max@sel And this is how to select the right combination. Note that we take the old maximum information only if the size does not change in either direction. If the size has grown, we of course have a new maximum. If the size has shrunk, the old maximum (which was explicitly not enforced because of being in the current chunk) is invalid, and we start with this chunk as the new size. Note that even in the case of equality we must use the \the\wd\tw@ construct instead of #2 because #2 might be read in from the file, and so could have \catcode 11 versions of p and t in it which we want to be replaced by their ‘proper’ \catcode 12 versions.

297 \def\LT@max@sel#1#2{%
298 {\ifdim#2=\wd\tw@
299  #1%
300 \else
301  \number\c@LT@chunks
302 \fi}%
303 {\the\wd\tw@}}

9.9 The \hline Command

\LT@hline \hline and \hline\hline both produce two lines. The only difference being the glue and penalties between them. This is so that a page break at a \hline produces a line on both pages.^{6} Also this \hline is more like a \cline{1-\LT@cols}. tabular’s \hline would draw lines the full width of the page.

304 \def\LT@hline{%
305 \noalign{\ifnum0=`}\fi
306  \penalty\@M
307  \futurelet\@let@token\LT@@hline}

\LT@@hline This code is based on \cline. Two copies of the line are produced, as described above.

308 \def\LT@@hline{%
309 \ifx\@let@token\hline
310  \global\let\@gtempa\@gobble
311  \gdef\LT@sep{\penalty-\@medpenalty\vskip\doublerulesep}%
312 \else
313  \global\let\@gtempa\@empty
314  \gdef\LT@sep{\penalty-\@lowpenalty\vskip-\arrayrulewidth}%
315 \fi
316 \ifnum0=`{\fi}%
317 \multispan\LT@cols
318  \unskip\leaders\hrule\@height\arrayrulewidth\hfill\cr
319 \noalign{\LT@sep}%
320 \multispan\LT@cols
321  \unskip\leaders\hrule\@height\arrayrulewidth\hfill\cr
322 \noalign{\penalty\@M}%
323 \@gtempa}

9.10 Captions

\LT@caption The caption is \multicolumn{\LT@cols}{c}{⟨a parbox with the table’s caption⟩}

324 \def\LT@caption{%
325 \noalign\bgroup
326  \@ifnextchar[{\egroup\LT@c@ption\@firstofone}\LT@capti@n}

\LT@c@ption Caption command (with [optional argument]). \protect added in Version 3. \fnum@table added at V3.05.

327 \def\LT@c@ption#1[#2]#3{%

6longtable has always done this, but perhaps it would be better if hlines were omitted at a page break, as the head and foot usually put a hline here anyway.

328 \LT@makecaption#1\fnum@table{#3}%
329 \def\@tempa{#2}%
330 \ifx\@tempa\@empty\else
331  {\let\\\space
332  \addcontentsline{lot}{table}{\protect\numberline{\thetable}{#2}}}%
333 \fi}

\LT@capti@n Caption command (no [optional argument])

334 \def\LT@capti@n{%
335 \@ifstar
336  {\egroup\LT@c@ption\@gobble[]}%
337  {\egroup\@xdblarg{\LT@c@ption\@firstofone}}}

\LT@makecaption Put the caption in a box of width 0pt, so that it never affects the column widths. Inside that is a \parbox of width \LTcapwidth.

338 \def\LT@makecaption#1#2#3{%
339 \LT@mcol\LT@cols c{\hbox to\z@{\hss\parbox[t]\LTcapwidth{%

Based on article class \@makecaption, #1 is \@gobble in star form, and

 \@firstofone otherwise.
340  \sbox\@tempboxa{#1{#2: }#3}%
341  \ifdim\wd\@tempboxa>\hsize
342  #1{#2: }#3%
343  \else
344  \hbox to\hsize{\hfil\box\@tempboxa\hfil}%
345  \fi
346  \endgraf\vskip\baselineskip}%
347 \hss}}}

9.11 The Output Routine

The method used here for interfacing a special purpose output routine to the standard L^{A}TEX routine is lifted straight out of F. Mittelbach’s multicol package.

\LT@output Actually this is not so bad, with FM leading the way.

348 \def\LT@output{%
349 \ifnum\outputpenalty <-\@Mi
350  \ifnum\outputpenalty > -\LT@end@pen

If this was a float or a marginpar we complain.

351  \LT@err{floats and marginpars not allowed in a longtable}\@ehc
352  \else

We have reached the end of the table, on the scroll at least,

353  \setbox\z@\vbox{\unvbox\@cclv}%
354  \ifdim \ht\LT@lastfoot>\ht\LT@foot

The last foot might not fit, so:^{7}

355  \dimen@\pagegoal
356  \advance\dimen@-\ht\LT@lastfoot
357  \ifdim\dimen@<\ht\z@
358  \setbox\@cclv\vbox{\unvbox\z@\copy\LT@foot\vss}%

7An alternative would be to vsplit off a bit of the last chunk, so that the last page did not just have head and foot sections, but it is hard to do this in a consistent manner.

359  \@makecol
360  \@outputpage
361  \setbox\z@\vbox{\box\LT@head}%
 End of \ifdim\dimen@<\ht\@cclc.
362  \fi
 End of \ifdim \ht\LT@lastfoot > \ht\LT@foot.
363  \fi
 Reset \@colroom.
364  \global\@colroom\@colht
365  \global\vsize\@colht

Put the last page of the table on to the main vertical list.

366  \vbox
367  {\unvbox\z@\box\ifvoid\LT@lastfoot\LT@foot\else\LT@lastfoot\fi}%
 End of \ifnum\outputpenalty > -\LT@end@pen.
368  \fi
 Else \outputpenalty > -\@Mi.
369 \else

If we have not reached the end of the table,

370  \setbox\@cclv\vbox{\unvbox\@cclv\copy\LT@foot\vss}%
371  \@makecol
372  \@outputpage
 Reset \vsize.
373  \global\vsize\@colroom

Put the head at the top of the next page.

374  \copy\LT@head
 End of \ifnum\outputpenalty <-\@Mi.
375 \fi}

9.12 Commands for the the table head and foot

\LT@end@hd@ft The core of \endhead and friends. Store the current chunk in the box specified by #1. Issue an error if the table has already started. Then start a new chunk.

376 \def\LT@end@hd@ft#1{%
377 \LT@echunk

Changed from \relax to \endgraf at V3.04, see \LT@start.

378 \ifx\LT@start\endgraf
379  \LT@err
380  {Longtable head or foot not at start of table}%
381  {Increase LTchunksize}%
382 \fi
383 \setbox#1\box\z@
384 \LT@get@widths
385 \LT@bchunk}
\endfirsthead \endhead \endfoot \endlastfoot Call \LT@end@hd@ft with the appropriate box.
386 \def\endfirsthead{\LT@end@hd@ft\LT@firsthead}
387 \def\endhead{\LT@end@hd@ft\LT@head}
388 \def\endfoot{\LT@end@hd@ft\LT@foot}
389 \def\endlastfoot{\LT@end@hd@ft\LT@lastfoot}

9.13 The \multicolumn command

Earlier versions needed a special ‘draft’ form of \multicolumn. This is not needed in version 4, and so these commands have been removed.

\LTmulticolumn
\LT@mcwarn

9.14 Footnotes

The standard \footnote command works in a c column, but we need to modify the definition in a p column to overcome the extra level of boxing. These macros are based on the array package, but should be OK for the standard tabular.

\LT@startpbox Add extra code to switch the definition of \@footnotetext.

390 \def\LT@startpbox#1{%
391 \bgroup
392  \let\@footnotetext\LT@p@ftntext
393  \setlength\hsize{#1}%
394  \@arrayparboxrestore
395  \vrule \@height \ht\@arstrutbox \@width \z@}

\LT@endpbox After the parbox is closed, expand \LT@p@ftn which will execute a series of

\footnotetext[⟨num⟩]{⟨note⟩}

commands. After being lifted out of the parbox, they can migrate on their own from here.

396 \def\LT@endpbox{%
397 \@finalstrut\@arstrutbox
398 \egroup
399 \the\LT@p@ftn
400 \global\LT@p@ftn{}%
401 \hfil}

\LT@p@ftntext Inside the ‘p’ column, just save up the footnote text in a token register.

402 \def\LT@p@ftntext#1{%
403 \edef\@tempa{\the\LT@p@ftn\noexpand\footnotetext[\the\c@footnote]}%
404 \global\LT@p@ftn\expandafter{\@tempa{#1}}}%

405 ⟨/package⟩